if f(a+b) = f(a) + f(b) for all real numbers. Show how f(2a) = 2*f(a) and show how f(0) = 0.
Found 2 solutions by HyperBrain, Doni305:
Answer by HyperBrain(694)

f(a+b) = f(a) + f(b)
Replace b with a:
f(a+a)=f(a)+f(a)
f(2a)=2*f(a)

Answer by Doni305(1)

The qoutient of a number and 20 is four tenths. What is the number?